The length of a rectangular boat storage shed must be 4 meters more than the width, and the perimeter must be at least 112 meters. What is the range of values for the width?
This is how I set it up.
4x = 112
x = 112/4
x = 28
The answer key answer is "at least 26 meters"
Where did I go wrong? Please show me where I make mistakes from this word problem.
Thank you
jpa

Answer by bmckenzie40(10)

Dear jpa:
You made the honest mistake of solving for a square instead of a rectangle.
If you make the width as X then the length has to be X+4 which would make the equation look like this
2x + 2(X+4) = 112 therefore
2X + 2X + 8 = 112 and then
4X + 8 = 112
4X = 104
X = 26
I hope this helps you